OpenStudy (anonymous):

find the value of x so that the line passing through (6,-25) and (x,20) has a slope of -3

OpenStudy (anonymous):

\[ \frac{20-(-25)}{x-6}=-3 \] Solve for x

OpenStudy (anonymous):

\[ 45= -3(x-6)\\ 15= -x+6\\ \] How much is x?

OpenStudy (anonymous):

x=-9

OpenStudy (anonymous):

x=-9

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

YW